ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65 Bacopoulos, Alexis, National Technical University of Athens On the link among vectorial minimal projections and best approximations We examine some common structure of vectorial minimal projections and best approximations. Computational results are also presented. Reference: A. Bacopoulos and B. L. Chalmers, Approximation Theory VIII, Chui and Schumaker (eds.), pp. 15 - 22. World Scientific Pub. Co., 1995.